OpenStudy (anonymous):

What polynomial has roots of -4, 1, and 6 a.x3 - 3x2 - 22x + 24 b.x3 - x2 - 26x - 24 c.x3 + x2 - 26x + 24 d.x3 + 3x2 + 14x - 24

OpenStudy (jdoe0001):

multiply the roots and expand -4 => x= -4 => x+4 = 0 => (x+4) as the root 1 => x = 1 => x-1 = 0 => (x-1) as the root ....

OpenStudy (anonymous):

similar to what i started writing

OpenStudy (anonymous):

@jdoe0001 can you explain a little bit further?

OpenStudy (anonymous):

@tkhunny that was kind of interesting, where did you learn that trick?

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

Be a little careful. In your example, the last term is (-a)(-b)(-c)(-d), and the negatives cancel out. Will that happen in our problem?

OpenStudy (tkhunny):

Be a LOT more careful. Not sure how I talked myself into that. Once we have the FACTORS from the ROOTS, THEN it's as obvious as I said. Root -4 1 6 Factor x - (-4) = x+4 x - 1 x - 6 Now, (+4)(-1)(-6) = +24 and immediately discard B and D. Yikes, time for a break.

OpenStudy (anonymous):

Basically, the fact that -4, 1, and 6 are roots of this polynomial tell us that:\[(x-(-4))(x-1)(x-6)\]is the polynomial we are after. Your job is to multiply this all together. That will give you the answer.

OpenStudy (anonymous):

Note this is the same as:\[(x+4)(x-1)(x-6)\]

OpenStudy (anonymous):

@joemath314159 alright, alright, I can understand what you saying :) thanks a lot!